### v1

# Social Optimization in Noncooperative Games under Central Regulation

Abstract
This paper proposes a novel optimization problem building on noncooperative games under central regulation, which can be formulated as a bilevel structure. In the low-level, each player competes to minimize its own cost function that depends not only on the strategies of all players, but also on an intervention decision of the central regulator, while the central regulator located at the high-level attempts to achieve the social optimum, that is, to minimize the sum of cost functions of all players through an adjustable intervention decision. In this setting, under the intervention of the central regulator, the low-level players perform in a noncooperative game and aim to seek the Nash equilibrium, which indeed is related with the regulator’s decision. Meanwhile, the objective of the regulator is to choose a decision such that the social cost, i.e., the sum of cost functions of all players is minimum. This formulated bilevel social optimization problem is proven to be constrained, nonconvex and nonsmooth. To address this intricate problem, an inexact zeroth-order algorithm is developed by virtue of the smoothing techniques, allowing for the Nash equilibrium of the low-level game to be computed in an inexact manner. Levering the properties of smoothing techniques, it is rigorously shown that the devised algorithm achieves a sublinear convergence rate for computing a stationary point of a related optimization problem with a smoothed objective. Moreover, the sublinear convergence rate in the scenario where the exact equilibrium of the low-level game is available is also discussed. Finally, numerical simulations are conducted to demonstrate the efficiency of theoretical findings.
Index Terms:
Social optimum, noncooperative games, nondifferentiable nonconvex optimization, smoothing techniques, gradient-free method.
I Introduction
Game theory provides a suitable framework for modeling strategic interactions among rational players that aim to selfishly minimize their private, yet interdependent objective functions. This competitive setting has been commonly applied to various domains, including network congestion control [1], demand-side management in smart grids [2], and social networks [3]. One of most important issues in noncooperative games is to seek Nash equilibria (NEs) or generalized Nash equilibria (GNEs) which eventually dictate stable and satisfactory states wherein no player possesses an incentive for unilateral deviation.
In the past decades, noncooperative games have been extensively explored under different scenarios. Generally speaking, the existing algorithms for seeking NEs of noncooperative games can be grouped into three categories. The simplest setup pertains to unconstrained games [4, 5], wherein each player possesses the liberty to select strategies from the entire real number or vector spaces. Moving towards a slightly more intricate case, games with local uncoupled decision set constraints are considered [6, 7], necessitating players to confine their decisions within predetermined nonempty sets, which are often presumed to be closed and convex. Furthermore, much attention is paid for games involving coupled equality or inequality constraints [8, 9].
It should be noted that in noncooperative games, each player acts selfishly and only concerns its individual interest. However, such self-centered behaviors may lead to deterioration in performance of the social cost, that is, the sum of cost functions of all players. Specially, some existing studies have yielded valuable insights into the detrimental consequences of selfish behaviors on performance degradation in diverse fields, such as supply chains [10], resource allocation [11], and congestion games [12, 13] by virtue of the metrics of the price of anarchy and the price of stability [14, 15]. A prospective direction of research involves improving social cost while also accommodating the preferences of individual players in noncooperative games. Consequently, a central regulator has its purpose to regulate its decision to eusure that the social cost is minimum. For instance, in the context of economic investment problems within network games, the central blank, who wishes to ensure market stability and avoid systemic risk, maximizes the overall social return for all participants by meticulously regulating individual standalone marginal return [16]. Similarly, in power systems, the total social resource consumption is minimized through the strategic adjustment of electricity prices [17].
To date, there has been some literature on adaptive incentive mechanisms for steering NEs towards a desirable point [18, 19, 20, 21, 22]. However, these works were explored from the control theoretic perspective, and paid attention to sufficient conditions or direct computation of the best targeting intervention for problems exhibiting specific structures. Recently, combining with the gradient dynamics, the authors in [16] studied continuous-time intervention protocols based on limited knowledge, which were rigorously shown to successfully drive actions of players in quadratic network games to the social optimum. For nonconvex cooperative aggregative games, a distributed stochastic annealing algorithm was developed in [23] without external intervention and was proved to weakly converge. Note that both of [16] and [23] have not yet derived any result regarding the convergence rate. Furthermore, most of the aforementioned works did not consider games with set constraints and intervention constraints. In [24], a bilevel incentive design algorithm is devised for probability simplex constrained games and the sublinear rate of convergence is provided.
It is emphasized that the exploration of social optima in noncooperative games is devoted to the global objective of the social optimization problem, thereby facilitating the central regulator in further improving the resource allocation efficiency. From a theoretical standpoint, when players’ objective functions hinge on their own decisions and those of other players, as well as a decision determined by the central regulator. Then, the corresponding situation extends beyond the scope of traditional noncooperative games, shifting the focus towards achieving a global optimal objective. This problem becomes increasingly complex and challenging due to its inherent bilevel structure, particularly, the mutual coupling of decisions between players and the central regulator. Therefore, the problem of achieving the social optimum in general noncooperative games remains to be an open and challenging issue.
Motivated by the above analysis, to overcome the obstacle that the behaviors of selfish players proposing to minimize their own costs may lead to that the social cost is not optimal, this paper investigates the social optimization problem, which inherently features a bilevel structure. In the low-level, players participate in a noncooperative game, where each individual cost function relies not only on all players’ strategies but also on an intervention decision of a central regulator. Under the intervention of the regulator, all players endeavor to seek the NE corresponding to the regulator’s decision. Then, in the high-level, the central regulator tries to find an optimal decision so as to achieve the social optimum. The significant challenges arise from the absence of a closed-form expression for the low-level game, as well as the inherent nondifferentiability, nonconvexity, and constraints of the high-level function. An inexact gradient-free method by utilizing smoothing techniques is devised to solve this problem and elaborately analyzed to exhibit sublinear convergence rate. The main contributions are summarized as follows.
  • i)
    A new social optimization framework based on noncooperative games is formulated for the first time which is inherently equipped with a bilevel structure, where the strategies of low-level players and the decisions of the high-level regulator are intricately and mutually coupled. Although the related problems regarding social optimization were investigated in [23, 16, 24], the problem settings are rather different. Specifically, there was no intervention involved in [23]. In fact, players are reluctant to sacrifice their personal wellbeing to move toward the social optimum without external intervention. Moreover, to design interventions for special quadratic network games [16], the regulator is presumed to be aware of the social optimum in advance. However, this is often difficult or even impossible in reality. Therefore, our concerned problem is more general and practical.
  • ii)
    The formulated problem is indeed constrained, nonsmooth, and nonconvex, leading to difficulties in problem solving. By employing the randomized smoothness and Moreau smoothness techniques, the studied problem is transformed into a corresponding smoothed unconstrained optimization problem. Then a zeroth-order algorithm is proposed, which enables the low-level game to be solved in an inexact manner to contend with the unavailable exact equilibrium of the low-level game. The devised algorithm utilizes random gradient-free oracles instead of relying on direct gradient information, and does not require the exact equilibrium of the low-level game. As a result, the algorithm becomes more practical and flexible in scenarios where obtaining gradients of the objective functions is infeasible or the exact NE is prohibitively expensive to evaluate.
  • iii)
    Through rigorous analysis, in the case where the central regulator cannot obtain the NE of the low-level game, the sublinear convergence rate is established for computing a stationary point of the corresponding smoothed unconstrained optimization problem within the inexact regime. Furthermore, the sublinear rate is also derived in the scenario where the exact NE of the low-level game is available. However, the convergence rate of algorithms for social optimization in [16, 23] was not established.
The rest of this paper is outlined as follows. Section II formulates the problem and introduces the preliminaries. A zeroth-order algorithm is provided in Section III, followed by the convergence results. In Section IV, numerical simulations are given to illustrate the theoretical findings. Section V makes a brief conclusion.
[…]
II PROBLEM FORMULATION AND PRELIMINARIES
In this section, the studied problem is first introduced and then some essential preliminaries related to smoothing techniques are presented.
II-A Problem Formulation
This paper introduces a novel social optimization framework grounded in noncooperative games consisting of $N$ players located at a low-level and a central regulator in the high-level. In the low-level, players perform in a noncooperative game and aim to selfishly minimize their own cost functions, dependent on not only the strategies of all players but also an adjustable intervention decisions of the central regulator. When the central regulator makes a decision, all players response at once simultaneously, aiming to seek the NE, which indeed is related to the regulator’s decision. Then the central regulator in the high-level seeks to determine an optimal intervention decision to reach the social optimum. Specifically, the schematic sketch of the problem setup is presented in Fig. 1, and the further expressions are detailed below.
Figure 1: Sketch of the problem setup. Each of the $N$ players in the low-level aims to minimize the personal cost function $f_{i}$ given intervention decision $\theta$ of the central regulator through a communication network. The regulator in the high-level utilizes required data received from all players to choose an optimal intervention decision to achieve social optimum. Here, $x_{-i}:=\operatorname*{col}\{x_{i}\}_{i\in[N]\backslash\{i\}}$ and $x(\theta):=\operatorname*{col}\{x_{i}(\theta)\}_{i\in[N]}$.
Low-level game. Each player $i\in[N]$ takes strategy $x_{i}$ from its strategy set $X_{i}\subseteq\mathbb{R}^{d_{i}}$. Let $x=\operatorname*{col}\{x_{i}\}_{i\in[N]}\in X:=X_{1}\times\cdots\times X_{N}\subseteq\mathbb{R}^{d}$ with $d=\sum_{i=1}^{N}d_{i}$ and $x_{-i}=\operatorname*{col}\{x_{i}\}_{i\in[N]\backslash\{i\}}\in{X_{-i}}:=X_{1}\times\cdots\times X_{i-1}\times X_{i+1}\times\cdots\times X_{N}\subseteq\mathbb{R}^{d-d_{i}}$ denote, respectively, the joint decision of all players and the joint strategy of all other players except $i$. The cost function of player $i$ is $f_{i}(x_{i},x_{-i},\theta)$ which depends on $x_{i}$, $x_{-i}$ and the central regulator’s intervention decision $\theta$ selected from a feasible set $\Theta\subseteq\mathbb{R}^{n}$. Herein, each player is not willing to share its own strategy and objective function with other players for the reason of privacy, and instead only has access to the information of partial players through local information exchange. Given a current decision $\theta$, players select their strategies with the aim of minimizing individual objective functions $f_{i}(x_{i},x_{-i},\theta)$, $i\in[N]$ to seek the NE that is in their best interest. This leads to a collection of interdependent optimization problems as follows:
$\min_{x_{i}\in X_{i}}f_{i}(x_{i},x_{-i},\theta),\forall i\in[N].$ (1)
In this study, we limit our focus to the case where the low-level game (1) admits a unique NE for any particular parametrization $\theta\in\Theta$. Upon completion of the game process, an NE $x(\theta)=(x(\theta),x_{-i}(\theta))$ is obtained, which corresponds to a set of cost functions of players: $f_{i}(x_{i}(\theta),x_{-i}(\theta),\theta),i\in[N]$.
Definition 1.
For any central regulator’s decision $\theta\in\mathbb{R}^{n}$, a strategy profile $x(\theta)=(x_{i}(\theta),x_{-i}(\theta))\in X$ is an NE for game (1) if for every player $i\in[N]$, it holds that
$f_{i}(x_{i}(\theta),x_{-i}(\theta),\theta)\leq f_{i}(x_{i},x_{-i}(\theta),\theta),\forall x_{i}\in X_{i}.$ (2)
High-level optimization. After receiving all players’ feedback $x_{i}(\theta)$, $i\in[N]$, the central regulator endeavors to determine the optimal decision $\theta^{*}$ that minimizes the sum of cost functions of all players, thereby achieving the social optimum. More precisely, the regulator aims to solve the following optimization problem:
$\min_{\theta\in\Theta}f(x(\theta),\theta):=\sum_{i=1}^{N}f_{i}(x(\theta),\theta).$ (3)
Note that the information on the players is available to the central regulator, making it preferable to leverage required data for effective coordination and decision-making.
In conclusion, the studied bilevel social optimization problem can be formulated as
$\min_{\theta\in\Theta}\sum_{i=1}^{N}f_{i}(x_{i}(\theta),x_{-i}(\theta),\theta),$
[…]
Consequently, the main objective of this paper is to seek the optimal decision $\theta^{*}$ to ultimately achieve the social optimum.
Remark 1.
Note that the key distinction between the social optimization (II-A) and Stackelberg games studied in [25, 26, 27] lies in the role of the high-level leader. Specifically, in the realm of Stackelberg games, the leader pursues the minimization of its own objective while adhering to the equilibrium constraints established by the low-level game among the followers. However, the central regulator in social optimization is tasked with optimizing the social cost. Moreover, in noncooperative games, only players compete with each other to minimize their own cost functions and no other regulators are involved. Therefore, the studied problem is essentially different from these frameworks since it is driven by the overall benefit to society, with a central regulator coordinating the whole network to reach this goal.
Some standard assumptions are listed next.
Assumption 1.
  • i)
    The set $X_{i}$ for any $i\in[N]$ is nonempty, compact, and convex.
  • ii)
    The set $\Theta$ is nonempty, closed, and convex.
  • iii)
    For any $i\in[N]$, $f_{i}(x_{i},x_{-i},\theta)$ is continuously differentiable and convex with respect to $x_{i}$ for any $x_{-i}\in\mathbb{R}^{d-d_{i}}$ and $\theta\in\mathbb{R}^{n}$.
  • iv)
    For any $i\in[N]$, $f_{i}(x,\theta)$ is $L_{\theta}$-Lipschitz continuous with respect to $\theta$ for any $x\in X$, i.e., for any $\theta,\theta^{\prime}\in\mathbb{R}^{n}$, there exists a constant $L_{\theta}>0$ such that
    $\|f_{i}(x,\theta)-f_{i}(x,\theta^{\prime})\|\leq L_{\theta}\|\theta-\theta^{\prime}\|.$ (5)
From Assumption 1-i) and iii), it can be further obtained that for any $i\in[N]$, $\theta\in\mathbb{R}^{n}$, and $x,x^{\prime}\in X$, there exist $B_{X},L_{x}>0$ such that
$\|x\|\leq B_{X},\ \|f_{i}(x,\theta)-f_{i}(x^{\prime},\theta)\|\leq L_{x}\|x-x^{\prime}\|.$ (6)
Given $\theta\in\mathbb{R}^{n}$, the pseudo-gradient mapping $G:\mathbb{R}^{d}\times\mathbb{R}^{n}\to\mathbb{R}^{d}$ is defined to be
$G(x,\theta)=\operatorname*{col}\{\nabla_{i}f_{i}(x_{i},x_{-i},\theta)\}_{i\in[N]}.$ (7)
Assumption 2.
  • i)
    The pseudo-gradient $G$ is $\mu$-strongly monotone ($\mu>0$) and $l$-Lipschitz continuous ($l>0$) with respect to $x$ for any $\theta\in\mathbb{R}^{n}$, i.e., for any $x,x^{\prime}\in\mathbb{R}^{d}$,
    $(G(x,\theta)-G(x^{\prime},\theta))^{\textit{T}}(x-x^{\prime})\geq\mu\|x-x^{\prime}\|^{2},$ (8)
    $\|G(x,\theta)-G(x^{\prime},\theta)\|\leq l\|x-x^{\prime}\|.$ (9)
  • ii)
    $G$ is $l_{\theta}$-Lipschitz continuous ($l_{\theta}>0$) with respect to $\theta$ for any $x\in X$, i.e., for any $\theta,\theta^{\prime}\in\mathbb{R}^{n}$,
    $\|G(x,\theta)-G(x,\theta^{\prime})\|\leq l_{\theta}\|\theta-\theta^{\prime}\|.$ (10)
Under Assumption 1-i) and iii), it follows from Proposition 1.4.2 in [28] that for any given $\theta\in\mathbb{R}^{n}$, a solution $x(\theta)\in X$ is an NE of game (1) if and only if
${\bf 0}\in G(x(\theta),\theta)+\mathrm{N}_{X}(x(\theta)).$ (11)
The existence of the NE is guaranteed from Corollary 2.2.5 in [28]. Additionally, Assumption 2-i) ensures the uniqueness of the NE by Theorem 2.3.3 in [28], which is also made in [7, 29, 4, 6, 9].
Lemma 1.
Under Assumptions 1 and 2, $f_{i}(x(\theta),\theta)$ is Lipschitz continuous, i.e., there exists a constant $L_{F}>0$ satisfying
$\|f_{i}(x(\theta),\theta)-f_{i}(x(\theta^{\prime}),\theta^{\prime})\|\leq L_{F}\|\theta-\theta^{\prime}\|,\forall\theta\in\mathbb{R}^{n}.$ (12)
Proof.
See Appendix A. ∎
The communication pattern among players is characterized by a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{A})$, where $\mathcal{V}=[N]$ is the player set, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the directed edge set, and $\mathcal{A}=[a_{ij}]\in\mathbb{R}^{N\times N}$ is the adjacency matrix. $a_{ij}>0$ if $(j,i)\in\mathcal{E}$ or $i=j$, and $a_{ij}=0$, otherwise. Here, a directed edge $(j,i)\in\mathcal{E}$ means player $j$ could send information to player $i$. Let $\mathcal{N}_{i}:=\{j\in[N]:(j,i)\in\mathcal{E}\}$ be the set of neighbors of player $j$. A directed path from $i_{1}$ to $i_{s}$ is defined as a sequence of edges $(i_{p},i_{p+1})\in\mathcal{E},p\in[s-1]$ for different nodes. A directed graph is said to be strongly connected if there is a directed path from every node to any other node. It is assumed that $a_{ii}>0$ for all $i\in[N]$ in this paper. In the following, a standard condition on players’ interaction is imposed.
Assumption 3.
$\mathcal{G}$ is strongly connected. The adjacency matrix $\mathcal{A}$ is doubly stochastic, i.e., $\mathcal{A}\bm{1}_{N}=\mathcal{A}^{\textit{T}}\bm{1}_{N}=\bm{1}_{N}$.
Under Assumption 3, there exists $\bar{\sigma}\in(0,1)$ such that $\sigma_{N-1}(\mathcal{A})\leq\bar{\sigma}$.
Generally speaking, by virtue of the properties of objective functions and constraint sets of the players, various methods can be employed to compute NEs of the low-level game (1). Hence, our main attention shifts towards addressing the high-level optimization problem (3). From the preceding analysis, it is clear that to minimize the high-level function $f(x(\theta),\theta)$ with respect to $\theta$, one needs to firstly solve the low-level game problem (1), while its corresponding optimal decision $x(\theta)$ in turn depends on $\theta$. This mutual coupling between $x$ and $\theta$ renders problem (II-A) intrinsically hard to solve. Naturally, a direct and efficient approach involves gradient-based methods, where an implicit gradient of the high-level function $f(x(\theta),\theta)$ is required to be computed by utilizing the chain rule, as follows
$\nabla f(x(\theta),\theta)$ $=\nabla x(\theta)\nabla_{x}f(x(\theta),\theta)+\nabla_{\theta}f(x(\theta),\theta).$
Subsequently, the high-level decision variable $\theta$ is updated utilizing the projected gradient descent method. However, several challenges arise:
  • i)
    For a particular $\theta\in\mathbb{R}^{n}$, obtaining a closed-form characterization of $x(\theta)$ is often computationally expensive or even unavailable unless the functions $f_{i}$, $i\in[N]$ take very special forms. Hence, this typically limits the applicability of standard first-order and zeroth-order schemes.
  • ii)
    The computation of implicit gradient $\nabla f(\theta,x(\theta))$ requires not only access to $x(\theta)$ but also the assumption of differentiability of $x(\theta)$. Unfortunately, as demonstrated by Example 1 below, $x(\theta)$ and hence $f(x(\theta),\theta)$ are non-differentiable in general, rendering the gradient-based methods infeasible. Furthermore, one may not compute subgradients or Clarke generalized gradients easily in such settings either [30]. As a consequence, the design and analysis of computational methods for addressing problem (II-A) are considerably challenging.
  • iii)
    The function $f(x(\theta),\theta)$ may exhibit nonconvexity in $\theta$, as demonstrated by Example 1 below. This imposes significant challenges in the convergence analysis and, especially, in deriving the convergence rate.
Example 1.
Consider a noncooperative game with $2$ players. The cost function of player $i$ is $f_{i}(x_{i},x_{-i},\theta)=x_{i}^{2}-2x_{-i}-2x_{i}\theta$ and the strategy set is $X_{i}=[2/3,1]$, $\forall i\in[2]$. The feasible set of the central regulator is $\Theta=[0,1]$. Based on the optimal condition (11), it can be derived that
$x_{i}(\theta)=\left\{\begin{aligned} &2/3,&\text{if}\ \theta\in[0,2/3],\\ &\theta,&\text{if}\ \theta\in[2/3,1].\end{aligned}\right.$
Through simple calculation, one has
$f(x(\theta),\theta)$ $=\sum_{i=1}^{2}f_{i}(x(\theta),\theta)=\left\{\begin{aligned} &-\frac{4}{3}\theta-\frac{8}{9},&\text{if}\ \theta\in[0,2/3],\\ &-2\theta^{2}-4\theta,&\text{if}\ \theta\in[2/3,1].\end{aligned}\right.$
Obviously, at the point $\theta=2/3$ the mapping $x(\theta)$ and thus $f(x(\theta),\theta)$ are not differentiable. In addition, when $\theta\in[2/3,1]$, $f(x(\theta),\theta)=-2\theta^{2}-4\theta$ is nonconvex.
II-B Randomized Smoothness
Consider a function $g:\mathbb{R}^{n}\to\mathbb{R}$. Define the uniform smoothing $\hat{g}$ of $g$ as
$\hat{g}(z):=\mathrm{E}_{\nu\in\mathbb{B}^{n}}[g(z+\xi\nu)],\ \forall z\in\mathbb{R}^{n},$ (13)
where $\xi>0$ is the smoothing parameter. Then a two-point sampling gradient estimator is proposed as [31]
^$\hat{\nabla}g(z)=\frac{n}{\xi}(g(z+\xi u)-g(z))u,$ (14)
where $u\in\mathbb{S}^{n}$ is a stochastic vector with uniform distribution.
The following lemma shows some properties of $\hat{g}$ and ^$\hat{\nabla}g$.
Lemma 2.
[31]
  • i)
    $\hat{g}(z)$ is differentiable even when $g$ is not, it holds that
    $\nabla\hat{g}(z)=\mathrm{E}_{u\in\mathbb{S}^{n}}[\hat{\nabla}g(z)].$ (15)
  • ii)
    If $g(z)$ is $L_{g}$-Lipschitz continuous ($L_{g}>0$), then $\nabla\hat{g}(z)$ is $\frac{nL_{g}}{\xi}$-Lipschitz continuous. Moreover, it holds that
    ^$\|\hat{\nabla}g(z)\|\leq nL_{g}.$ (16)
II-C Moreau Smoothness
Let $\chi_{\Theta}:\mathbb{R}^{n}\to\mathbb{R}\cup+\infty$ be an indicator function defined on a nonempty closed and convex set $\Theta\subseteq\mathbb{R}^{n}$, that is,
$\chi_{\Theta}(\theta):=\left\{\begin{aligned} &0,&\text{if}\ \theta\in\Theta,\\ &+\infty,&\text{if}\ \theta\notin\Theta.\end{aligned}\right.$
The Moreau smoothing $\hat{\chi}_{\Theta}$ of $\chi_{\Theta}$ is defined as [32]
$\hat{\chi}_{\Theta}(\theta):=\frac{1}{2\xi}{\operatorname*{dist}}^{2}(\theta,\Theta),$ (17)
where $\xi$ is the smoothing parameter. By Theorem 6.60 in [32], $\hat{\chi}_{\Theta}$ is differentiable and
$\nabla\hat{\chi}_{\Theta}(\theta)=\frac{1}{\xi}(\theta-\Pi_{\Theta}(\theta)).$
Moreover, $\nabla\hat{\chi}_{\Theta}(\theta)$ is $\frac{1}{\xi}$-Lipschitz continuous.
Remark 2.
Intuitively, the randomized smoothed version $\hat{g}$ and Moreau smoothed variant $\hat{\chi}_{\Theta}$, as substitutes for original functions $g$ and $\chi_{\Theta}$, respectively, are typically introduced to contend with the nonsmoothness, which greatly help develop a zeroth-order algorithm.
III MAIN RESULTS
In this section, a zeroth-order algorithm for addressing the constrained, nonsmooth, and nonconvex problem (II-A) is developed, which allows for utilizing inexact solutions of the low-level game. Moreover, the convergence rate of the proposed algorithm are rigorously derived.
For notational convenience, define $F_{i}(\theta):=f_{i}(x(\theta),\theta))$ and $F(\theta):=\sum_{i=1}^{N}F_{i}(\theta)$. Note that the constrained optimization problem (3) is equivalent to
$\min_{\theta\in\mathbb{R}^{n}}\mathbf{F}(\theta):=F(\theta)+\chi_{\Theta}(\theta).$ (18)
Consider a smoothed unconstrained approximation of (3) based on (13) and (17), given as
$\min_{\theta\in\mathbb{R}^{n}}\hat{\mathbf{F}}(\theta):=\hat{F}(\theta)+\hat{\chi}_{\Theta}(\theta).$ (19)
where $\hat{F}(\theta):=\sum_{i=1}^{N}\hat{F}_{i}(\theta)$ with $\hat{F}_{i}(\theta):=\mathrm{E}_{\nu\in\mathbb{B}^{n}}[F_{i}(\theta+\xi\nu)]$ and $\hat{\chi}_{\Theta}(\theta):=\frac{1}{2\xi}{\operatorname*{dist}}^{2}(\theta,\Theta)$ with $\xi>0$ being the smoothing parameter. Denote by $\hat{\mathbf{F}}^{*}:=\min_{\theta\in\mathbb{R}^{n}}\hat{\mathbf{F}}(\theta)$ the optimal value of problem (19).
Before proceeding, we formalize the relationship between the original nonsmooth problem (II-A) and its smoothed counterpart (19). Toward this end, the $\delta$-Clarke generalized gradient of $\mathbf{F}$ at $\theta$ is introduced [33], denoted by $\partial_{\delta}\mathbf{F}(\theta)$, as follows
$\partial_{\delta}\mathbf{F}(\theta):=\operatorname{conv}\{\zeta\ |\ \zeta\in\partial\mathbf{F}(\theta^{\prime}),\|\theta^{\prime}-\theta\|\leq\delta\}.$ (20)
Basically, the set $\partial_{\delta}\mathbf{F}(\theta)$ is shown to be nonempty, compact, and convex [33]. Inspired by Proposition 2 in [30], the following lemma is shown.
Lemma 3.
Consider problem (II-A) and let Assumptions 1 and 2 be satisfied. Then for any $\delta>0$, if $\nabla\hat{\mathbf{F}}(\theta)=\bm{0}$ and $\xi\leq\frac{\delta}{\max\{2,nNL_{F}\}}$, there hold $\bm{0}\in\partial_{\delta}\mathbf{F}(\theta)$ and $\|\theta-\Pi_{\Theta}(\theta)\|\leq\xi nNL_{F}$.
According to Lemma 3, any stationary point of the $\xi$-smoothed problem satisfies an approximate stationary property for the original problem. Furthermore, the distance between the stationary point and the feasible set $\Theta$ is $\mathcal{O}(\xi)$. These findings lay the foundation for developing effective schemes to compute approximate stationary points of minimization of $\mathbf{F}$ in nonconvex and nonsmooth regimes.
In order to deal with the smoothed problem (19), by virtue of Lemma 2, for every $i\in[N]$, two point sampling gradient approximation of $\hat{F}_{i}(\theta)$ is
^$\hat{\nabla}F_{i}(\theta):=\frac{n}{\xi}(f_{i}(x(\theta+\xi u),\theta+\xi u)-f_{i}(x(\theta),\theta))u.$
Hence
^$\hat{\nabla}F(\theta)=\frac{n}{\xi}\sum_{i=1}^{N}(f_{i}(x(\theta+\xi u),\theta+\xi u)-f_{i}(x(\theta),\theta))u.$ (21)
Based on the preceding discussion, it is impractical to obtain the exact solution $x(\theta)$. We carefully address this challenge by introducing inexact evaluation of $x(\theta)$, where players are allowed to compute an $\varepsilon$-approximation solution $x_{\varepsilon}(\theta)$ satisfying
$\mathrm{E}[\|x_{\varepsilon}(\theta)-x(\theta)\|^{2}|\theta]\leq\varepsilon.$ (22)
As a consequence, one needs to consider the inexact zeroth-order gradient, denoted by
^$\hat{\nabla}_{\varepsilon}F_{i}(\theta)=\frac{n}{\xi}(f_{i}(x_{\varepsilon}(\theta+\xi u),\theta+\xi u)-f_{i}(x_{\varepsilon}(\theta),\theta))u.$
Input: Stepsize $\alpha>0$, smoothing parameter $\xi>0$, and initial vector $\theta_{0}\in\Theta$.
1 for $k=0,1,\ldots,K-1$ do
2      Select vector $u_{k}\in\mathbb{S}^{n}$ independently and uniformly at random.
3      Send $\theta_{k}+\xi u_{k}$ and $\theta_{k}$ to all players.
4      Apply Algorithm 2 to receive inexact evaluations $x_{\varepsilon_{k}}(\theta_{k}+\xi u_{k})$ and $x_{\varepsilon_{k}}(\theta_{k})$.
5      Evaluate the inexact zeroth-order gradient approximation ^$\hat{\nabla}_{\varepsilon_{k}}F(\theta_{k})$ by (23).
6      Update $\theta_{k}$ as (24).
7
8      end for
Algorithm 1 Inexact Smoothing-Enabled Zeroth-Order Method for Social Optimization
In a nutshell, an inexact smoothing-enabled zeroth-order method is developed in Algorithm 1, while an inexact solution of $x(\theta)$ is computed by a fully distributed NE seeking algorithm, outlined in Algorithm 2, which was inspired by [7]. At each iteration $k$, firstly, the central regulator select vector $u_{k}\in\mathbb{S}^{n}$ independently and uniformly at random and then send $\theta_{k}+\xi u_{k}$ and $\theta_{k}$ to all players. Simultaneously, the central regulator makes an inexact call to Algorithm 2 twice to compute $x_{\varepsilon_{k}}(\theta_{k}+\xi u_{k})$ and $x_{\varepsilon_{k}}(\theta_{k})$. Specifically, at time $t$, each player $i$ maintains variables $x_{i,t}$ and ${\bf x}_{i,t}^{j}$ to represent its own decision and estimate of player $j$’s decision, respectively. Define ${\bf x}_{i,t}^{i}:=x_{i,t}$ and also ${\bf x}_{j,t}^{-i}:=\operatorname*{col}\{{\bf x}_{j,t}^{l}\}_{l\in[N]\setminus i}$. Let ${\bf x}_{i,t}:=\operatorname*{col}\{{\bf x}_{i,t}^{j}\}_{j\in[N]}\in\mathbb{R}^{d}$ and ${\bf x}_{t}:=\operatorname*{col}\{{\bf x}_{i}\}_{i\in[N]}\in\mathbb{R}^{Nd}$. In the partial-decision information setting, players update their estimates based on the information received from their neighbors by applying dynamic consensus technique (25a), and subsequently adjust their own decisions through a gradient descent step (25b). Herein, it is mentioned that the partial gradient $\nabla_{i}f_{i}$ is evaluated on the local estimate $\hat{{\bf x}}_{i,t}$, while not employing the actual decision $x_{i,t}$. After receiving the estimated feedback $x_{\varepsilon_{k}}(\tilde{\theta}_{k}):=\operatorname*{col}\{x_{i,t_{k}}\}_{i\in[N]}$ from the players, the inexact zeroth-order gradient ^$\hat{\nabla}_{\varepsilon_{k}}F(\theta_{k})$ is computed as
^$\hat{\nabla}_{\varepsilon_{k}}F(\theta_{k})=\sum_{i=1}^{N}\frac{n}{\xi}(f_{i}(x_{\varepsilon_{k}}(\theta_{k}+\xi u_{k}),\theta_{k}+\xi u_{k})-f_{i}(x_{\varepsilon_{k}}(\theta_{k}),\theta_{k}))u_{k}.$ (23)
At last, the central regulator updates its decision using gradient descent
$\theta_{k+1}=\theta_{k}-\alpha(\hat{\nabla}_{\varepsilon_{k}}F(\theta_{k})+\frac{1}{\xi}(\theta_{k}-\Pi_{\Theta}(\theta_{k}))),$ (24)
where $\alpha>0$ is the step size to be determined.
Input: Stepsize $\gamma>0$, and initial vector $x_{i,0}\in X_{i}$, ${\bf x}_{i,0}^{-i}\in\mathbb{R}^{d-d_{i}}$.
1 for $t=0,1,\ldots,t_{k}-1$ do
2      for $i=1,...,N$ in parallel do
3           Receive $\tilde{\theta}_{k}$ (that is, $\theta_{k}+\xi u_{k}$ or $\theta_{k}$) from the central regulator.
4           Update
$\hat{{\bf x}}_{i,t}$ $=\sum_{j=1}^{N}a_{ij}{\bf x}_{j,t},$ (25a)
$x_{i,t+1}$ $=\Pi_{X_{i}}[\hat{{\bf x}}_{i,t}^{i}-\gamma\nabla_{i}f_{i}(\hat{{\bf x}}_{i,t},\tilde{\theta}_{k})],$ (25b)
${\bf x}_{i,t+1}^{-i}$ $=\hat{{\bf x}}_{i,t}^{-i}.$ (25c)
5           end for
6
7           end for
Players send $x_{i,t_{k}},i\in[N]$ to the central regulator.
Algorithm 2 Fully Distributed NE Seeking
Remark 3.
Utilizing inexact estimation of the NE of the low-level game inherently introduces a bias in the approximation of the zeroth-order gradient. This bias, if uncontrolled, may undesirably propagate and impact the overall performance of the algorithm. Notably, the degree of inexactness in $x_{\varepsilon_{k}}(\tilde{\theta}_{k})$ is precisely controlled by the termination criterion $t_{k}$ in Algorithm 2. This criterion will be formally derived in subsequent Theorem 1 to establish the convergence rate. Additionally, note that the central regulator in the high-level does not need to acquire all information on the players. In fact, at moment $k+1$, only two function values $f_{i}(x_{\varepsilon_{k}}(\theta_{k}+\xi u_{k}),\theta_{k}+\xi u_{k})$ and $f_{i}(x_{\varepsilon_{k}}(\theta_{k}),\theta_{k})$ are required to be disclosed by the central regulator to update its decision $\theta_{k}$. In addition, it is noteworthy that in the setting of partial-decision information regarding the underlying game, each individual strategy and gradient of objective function are not accessible to other players. Consequently, every player must exert extra effort to estimate strategies of other players and hence steps (25a) and (25c) are introduced in Algorithm 2. However, during the practical implementation of Algorithm 1, it is only necessary to employ simple gradient descent method to update the players’ decision variables to obtain approximate solution $x_{\varepsilon_{k}}(\tilde{\theta}_{k})$ of the low-level game, which will simplify Algorithm 1 greatly.
Given $\theta\in\mathbb{R}^{n}$, define the extended pseudo-gradient mapping $\mathbf{G}:\mathbb{R}^{Nd}\times\Theta\to\mathbb{R}^{d}$ as
$\mathbf{G}({\bf x},\theta):=\operatorname*{col}\{\nabla_{i}f_{i}(x_{i},{\bf x}_{i}^{-i},\theta)\}_{i\in[N]}.$
Lemma 4.
[7, Lemma 1] Under Assumptions 1-iii) and 2-i), for any $\theta\in\mathbb{R}^{n}$, the mapping $\mathbf{G}$ is $l^{\prime}$-Lipschitz continuous with respect to ${\bf x}$ for some $\mu\leq l^{\prime}\leq l$, that is,
$\mathbf{G}({\bf x},\theta)-\mathbf{G}({\bf x}^{\prime},\theta)\leq l^{\prime}\|{\bf x}-{\bf x}^{\prime}\|,\ \forall{\bf x},{\bf x}^{\prime}\in\mathbb{R}^{d}.$
To proceed further analysis, the convergence result of Algorithm 2 for resolving the low-level game (1) is presented. The proof can be obtained following Theorem 1 in [7].
Lemma 5.
Suppose that Assumptions 1-i), iii), 2-i), and 3 hold. Let $\{x_{t}\}_{t=0}^{t_{k}}$ be the sequence generated by Algorithm 2 at the $k$-th iteration of Algorithm 1. If the stepsize $\gamma>0$ is chosen such that
$\gamma$ $<\min\{1,\frac{\bar{\sigma}}{3l},\frac{2\mu}{l^{2}},\frac{2\mu(1-\bar{\sigma}^{2})}{a}\},$ (26)
with $a:=\bar{\sigma}^{2}(2ll^{\prime}+l^{\prime 2}+4\mu l^{\prime}+2l^{2})+2(l^{2}l^{\prime 2}+\mu l^{\prime 2}+2l^{2}l^{\prime 2})\bar{\sigma}^{2}+2l^{2}l^{\prime 2}\bar{\sigma}^{2}$, then $x_{t}$ converges to $x(\tilde{\theta}_{k})$ at a linear rate. Specifically,
$\|{\bf x}_{t_{k}}-{\bf 1}_{N}\otimes x(\tilde{\theta}_{k})\|^{2}\leq q^{t_{k}}\|{\bf x}_{0}-{\bf 1}_{N}\otimes x(\tilde{\theta}_{k})\|^{2},$ (27)
where $q:=\|Q_{\gamma}\|\in(0,1)$ with
$Q_{\gamma}=\left[\begin{array}[]{ccc}1-\frac{2\gamma\mu}{N}+\frac{\gamma^{2}l^{2}}{N}&\frac{\gamma(l+l^{\prime})+\gamma^{2}ll^{\prime}}{\sqrt{N}}\bar{\sigma}\\ \frac{\gamma(l+l^{\prime})+\gamma^{2}ll^{\prime}}{\sqrt{N}}\bar{\sigma}&(1+2\gamma l+\gamma^{2}l^{2})\bar{\sigma}^{2}\end{array}\right].$
It follows from (27) in Lemma 5 that
$\varepsilon_{k}$ $:=\|x_{t_{k}}-x(\tilde{\theta}_{k})\|^{2}$
$\leq q^{t_{k}}\|{\bf x}_{0}-{\bf 1}_{N}\otimes x(\tilde{\theta}_{k})\|^{2}$
$\leq 2(\|{\bf x}_{0}\|^{2}+NB_{X}^{2})q^{t_{k}},$ (28)
where the last inequality is derived owing to (6).
Define the natural filtration
$\mathscr{F}_{0}:=\{\emptyset,\Omega\},\ \mathscr{F}_{k}:=\sigma\{\theta_{0},u_{t}:0\leq t\leq k-1\},\ \forall k\geq 1.$
It can be seen that $\theta_{k}$ generated by Algorithm 1 is adapted to $\mathscr{F}_{k}$ and $u_{k}$ is independent of $\mathscr{F}_{k}$. Then, the following lemma on properties of the zeroth-order gradient is derived.
Lemma 6.
Under Assumptions 1 and 2, there hold
$\mathrm{E}[\hat{\nabla}F(\theta_{k})+\nabla\hat{\chi}_{\Theta}(\theta_{k})|\mathscr{F}_{k}]=\nabla\hat{\mathbf{F}}(\theta_{k}),$ (29)
$\mathrm{E}[\|\hat{\nabla}_{\varepsilon_{k}}F(\theta_{k})-\hat{\nabla}F(\theta_{k})\|^{2}|\mathscr{F}_{k}]\leq 4N^{2}\frac{L_{x}^{2}n^{2}}{\xi^{2}}\varepsilon_{k},$ (30)
$\mathrm{E}[\|\hat{\nabla}F(\theta_{k})+\nabla\hat{\chi}_{\Theta}(\theta_{k})-\nabla\hat{\mathbf{F}}(\theta_{k})\|^{2}|\mathscr{F}_{k}]\leq N^{2}n^{2}L_{F}^{2}.$ (31)
Proof.
See Appendix B. ∎
Equipped with the previous preparations, it is ready to present the main result of this paper now.
Theorem 1.
Suppose that Assumptions 1-3 hold. Let $\{\theta_{k}\}$ be the sequence generated by Algorithm 1, if $\alpha\leq\frac{\xi}{4(nNL_{F}+1)}$, then it holds
$\frac{1}{K}\sum_{k=0}^{K-1}\mathrm{E}[\|\nabla\hat{\mathbf{F}}(\theta_{k})\|^{2}]$ $\leq\frac{4(\hat{\mathbf{F}}(\theta_{0})-\hat{\mathbf{F}}^{*})}{\alpha K}+8N^{2}n^{2}L_{F}^{2}\alpha\frac{nNL_{F}+1}{\xi}$
$\quad+\frac{20N^{2}L_{x}^{2}n^{2}\sum_{k=0}^{K-1}\varepsilon_{k}}{\xi^{2}K}.$ (32)
Proof.
Note that $F$ is $NL_{F}$-Lipschitz continuous, then one has from Lemma 2 that $\nabla\hat{F}$ is $\frac{nNL_{F}}{\xi}$-Lipschitz continuous. Moreover, $\nabla\hat{\chi}_{\Theta}$ is $\frac{1}{\xi}$-Lipschitz continuous. Hence, $\nabla\hat{\mathbf{F}}$ is $\frac{nNL_{F}+1}{\xi}$-Lipschitz continuous. In view of the update of $\theta_{k}$ in (24), it follows from Lemma 5.7 in [32] that
$\hat{\mathbf{F}}(\theta_{k+1})$ $=\hat{\mathbf{F}}(\theta_{k})-\alpha\nabla\hat{\mathbf{F}}(\theta_{k})^{\textit{T}}(\hat{\nabla}F(\theta_{k})+\nabla\hat{\chi}_{\Theta}(\theta_{k}))$
$\quad-\alpha\nabla\hat{\mathbf{F}}(\theta_{k})^{\textit{T}}(\hat{\nabla}_{\varepsilon_{k}}F(\theta_{k})-\hat{\nabla}F(\theta_{k}))$
$\quad+\frac{nNL_{F}+1}{2\xi}\alpha^{2}\|\hat{\nabla}F(\theta_{k})+\nabla\hat{\chi}_{\Theta}(\theta_{k})$
+^$\quad+\hat{\nabla}_{\varepsilon_{k}}F(\theta_{k})-\hat{\nabla}F(\theta_{k})\|^{2}$
$\leq\hat{\mathbf{F}}(\theta_{k})-\alpha\nabla\hat{\mathbf{F}}(\theta_{k})^{\textit{T}}(\hat{\nabla}F(\theta_{k})+\nabla\hat{\chi}_{\Theta}(\theta_{k}))$
$\quad+\frac{\alpha}{4}\|\nabla\hat{\mathbf{F}}(\theta_{k})\|^{2}+\alpha\|\hat{\nabla}_{\varepsilon_{k}}F(\theta_{k})-\hat{\nabla}F(\theta_{k})\|^{2}$
$\quad+\frac{nNL_{F}+1}{\xi}\alpha^{2}\|\hat{\nabla}F(\theta_{k})+\nabla\hat{\chi}_{\Theta}(\theta_{k})\|^{2}$
$\quad+\frac{nNL_{F}+1}{\xi}\alpha^{2}\|\hat{\nabla}_{\varepsilon_{k}}F(\theta_{k})-\hat{\nabla}F(\theta_{k})\|^{2},$
where the Young’s inequality $a_{1}^{T}a_{2}\leq\frac{\iota}{2}\|a_{1}\|^{2}+\frac{1}{2\iota}\|a_{2}\|^{2}$ for $\iota=2$ and relation $\|a_{1}+a_{2}\|^{2}\leq 2\|a_{1}\|^{2}+2\|a_{2}\|^{2}$ have been applied in the last inequality. Taking conditional expectation on both sides of the above inequality with respect to $\mathscr{F}_{k}$ and exploiting (29) in Lemma 6, one can obtain
$\mathrm{E}[\hat{\mathbf{F}}(\theta_{k+1})|\mathscr{F}_{k}]$
$\leq\hat{\mathbf{F}}(\theta_{k})-\frac{3\alpha}{4}\|\nabla\hat{\mathbf{F}}(\theta_{k})\|^{2}+2\frac{nNL_{F}+1}{\xi}\alpha^{2}\|\nabla\hat{\mathbf{F}}(\theta_{k})\|^{2}$
$\quad+2\frac{nNL_{F}+1}{\xi}\alpha^{2}\mathrm{E}[\|\hat{\nabla}F(\theta_{k})+\nabla\hat{\chi}_{\Theta}(\theta_{k})-\nabla\hat{\mathbf{F}}(\theta_{k})\|^{2}|\mathscr{F}_{k}]$
$\quad+(\frac{nNL_{F}+1}{\xi}\alpha^{2}+\alpha)\mathrm{E}[\|\hat{\nabla}_{\varepsilon_{k}}F(\theta_{k})-\hat{\nabla}F(\theta_{k})\|^{2}|\mathscr{F}_{k}].$
Taking expectation on both sides of the aforementioned inequality, then substituting (31) and (30) in Lemma 6 into it, and invoking $\alpha\leq\frac{\xi}{4(nNL_{F}+1)}$, it derives that
$\mathrm{E}[\hat{\mathbf{F}}(\theta_{k+1})]$ $\leq\mathrm{E}[\hat{\mathbf{F}}(\theta_{k})]-\frac{1}{4}\alpha\mathrm{E}[\|\nabla\hat{\mathbf{F}}(\theta_{k})\|^{2}]$
$+2N^{2}n^{2}L_{F}^{2}\frac{nNL_{F}+1}{\xi}\alpha^{2}+\frac{5N^{2}L_{x}^{2}n^{2}\alpha}{\xi^{2}}\varepsilon_{k}.$
Taking sum of the above inequality over $k$ from 00 to $K-1$ and dividing its both sides by $K$, it yields
$\frac{1}{K}\sum_{k=0}^{K-1}\mathrm{E}[\|\nabla\hat{\mathbf{F}}(\theta_{k})\|^{2}]$ $\leq\frac{4(\hat{\mathbf{F}}(\theta_{0})-\mathrm{E}[\hat{\mathbf{F}}(\theta_{K})])}{\alpha K}+8N^{2}n^{2}L_{F}^{2}\alpha\frac{nNL_{F}+1}{\xi}$
$\quad+\frac{20N^{2}L_{x}^{2}n^{2}\sum_{k=0}^{K-1}\varepsilon_{k}}{\xi^{2}K},$
which incorporating with the definition of $\hat{\mathbf{F}}^{*}$ completes the proof. ∎
Corollary 1.
Under the same conditions as in Theorem 1, when $t_{k}$ in Algorithm 2 is chosen such that $t_{k}\geq\lceil\frac{-s\ln(k+1)}{\ln q}\rceil$ with $s\in[0,1)$ and $\alpha=\frac{\alpha_{0}}{\sqrt{K}}$ with $\alpha_{0}>0$, there holds
$\frac{1}{K}\sum_{k=0}^{K-1}\mathrm{E}[\|\nabla\hat{\mathbf{F}}(\theta_{k})\|^{2}]=\mathcal{O}(\frac{1}{\sqrt{K}}+\frac{1}{K^{s}}).$ (33)
Proof.
Note that for any $s\in[0,1)$, there holds $\sum_{k=0}^{K-1}\frac{1}{(k+1)^{s}}\leq\frac{K^{1-s}}{1-s}.$ From $t_{k}\geq\lceil\frac{-s\ln(k+1)}{\ln q}\rceil$, one can derive $q^{t_{k}}\leq\frac{1}{(k+1)^{s}}$. Hence, it follows from (III) that $\sum_{k=0}^{K-1}\varepsilon_{k}\leq 2(\|{\bf x}_{0}\|^{2}+NB_{X}^{2})\frac{K^{1-s}}{1-s}.$ Then substituting the above inequality into (1) and noting the selection of $\alpha$ immediately ends the proof. ∎
Corollary 2.
Under the same conditions as in Theorem 1, when Algorithm 1 is executed in an exact way, i.e., ^$\hat{\nabla}_{\varepsilon_{k}}F(\theta_{k})$ is replaced by ^$\hat{\nabla}F(\theta_{k})$ defined in (21) in step $5$, and $\alpha\leq\frac{\xi}{2(nNL_{F}+1)}$, then
$\frac{1}{K}\sum_{k=0}^{K-1}\mathrm{E}[\|\nabla\hat{\mathbf{F}}(\theta_{k})\|^{2}]\leq\frac{2(\hat{\mathbf{F}}(\theta_{0})-\hat{\mathbf{F}}^{*})}{\alpha K}+2N^{2}n^{2}L_{F}^{2}\frac{nNL_{F}+1}{\xi}\alpha.$
Proof.
See Appendix C. ∎
Corollary 3.
Under the same conditions as in corollary 2, let $\alpha=\frac{\alpha_{0}}{\sqrt{K}}$ with $\alpha_{0}>0$, then it holds
$\frac{1}{K}\sum_{k=0}^{K-1}\mathrm{E}[\|\nabla\hat{\mathbf{F}}(\theta_{k})\|^{2}]=\mathcal{O}(\frac{1}{\sqrt{K}}).$
Remark 4.
From Corollary 1, it is evident that in the inexact regime, with an appropriately selected inner iteration number $t_{k}$, our proposed algorithm can attain a sublinear convergence rate of $\mathcal{O}(\frac{1}{\sqrt{K}}+\frac{1}{K^{s}})$ with $s\in[0,1)$. On the other hand, one can see from Lemma 5 that an increasing number of iteration $t_{k}$ derives a more accurate inner loop solution, which ultimately results in a faster convergence speed owing to a larger $s$. Specially, the order of magnitude of (33) stands at $\mathcal{O}(\frac{1}{\sqrt{K}})$ when choosing $t_{k}\geq\lceil\frac{-s\ln(k+1)}{\ln q}\rceil$ and $s\in[\frac{1}{2},1)$. Moreover, according to Corollary 3, when the low-level game can be solved exactly, i.e., $\varepsilon_{k}=0$$\forall k=0,\ldots,K-1$, the convergence rate is also $\mathcal{O}(\frac{1}{\sqrt{K}})$. Note that the bound $\mathcal{O}(\frac{1}{\sqrt{K}})$ recovers the existing convergence rate of nonconvex nonsmooth optimization in [34]. Additionally, it should be emphasized that by utilizing the continuity of $x(\theta)$, if $\theta_{k}$ converges to the optimal solution $\theta^{*}\in\hat{\Theta}^{*}$, then $x(\theta)$ will correspondingly converges to $x(\theta^{*})$.
IV NUMERICAL SIMULATION
Consider the problem of flexible electric vehicle charging control [35]. Specifically, there are $N$ residents in a neighborhood who commute daily using electric vehicles and charge their vehicles at a public charging station. The charging process is regulated by a central authority (such as a power company or government) through adjustable price policies $\theta$ within a specified range $\Theta=[\theta_{\min},\theta_{\max}]$. Due to battery capacity limitations, every resident $i$ has its charging consumption $x_{i}\in X_{i}=[0,x_{i}^{\max}]$ and aims to minimize its electricity bill $f_{i}(x_{i},x_{-i},\theta):=c_{i}(x_{i}-d_{i})^{2}+a_{i}x_{i}+\lambda_{i}(x_{i}-\sum_{i=1}^{N}x_{i}/N)^{2}+rx_{i}\theta$. Herein, $c_{i},d_{i},a_{i},r>0$ are constants, and $\lambda_{i}$ captures resident $i$’s sensitivity to deviation from average consumption. Moreover, $x_{i}\theta$ represents the charging cost imposed by the central regulator when resident $i$ selects a specific $x_{i}$. In order to maintain a balanced demand on the power grid demand and reduce total consumption of social resources, the central authority interacts with residents and is interested in miniming the overall cost of all residents.
Set $N=10$. For every $i\in[10]$, set $c_{i}=4+i$, $d_{i}=7+2i$, $a_{i}=10+i$, $\lambda_{i}=0.1$, and $r=1$. Moreover, let set constraints $X_{i}=[0,25]^{10}$ and $\Theta=[1,3]$. Choose stepsizes $\alpha=0.00001$ and $\gamma=0.01$, smoothing parameter $\xi=0.0001$, and inner iteration number $t_{k}=\lceil 5\ln(k+1)\rceil$. The evolution of residual $|\theta_{k}-\theta^{*}|$ versus iteration $k$ is plotted in Fig. 2, demonstrating that the regulator can find the optimal decision successfully. Furthermore, Fig. 3 gives rises to the trajectory of $|\nabla\hat{\mathbf{F}}(\theta_{k})|$, from which it can be seen that the proposed algorithm achieves a sublinear convergence rate.
Figure 2: The trajectory of $|\theta_{k}-\theta^{*}|$.
Figure 3: The trajectory of $|\nabla\hat{\mathbf{F}}(\theta_{k})|$.
V CONCLUSION
In this work, a new framework inspired by noncooperative games, i.e., social optimization, has been explored, in which each cost function depends not only on all players’ strategies, but also on an intervention decision of a central regulator. This problem possesses a bilevel structure, where the low-level players aim to seek the NE associated with the regulator’s decision in a noncooperative game, and the goal of the high-level regulator is to minimize the sum of all players’ objectives to achieve the social optimum. To address this constrained, nonconvex and nonsmooth problem, a gradient-free method was developed by virtue of the randomized smoothness and Moreau smoothness techniques. It was proved that the designed algorithm converges at a sublinear rate, irrespective of whether the underlying game is solved exactly or inexactly. Potential avenues for future research may be on stochastic objective functions and incorporate equality and inequality constraints.
Appendix A Proof of Lemma 1
Proof.
First, prove the Lipschitz continuity of $x(\theta)$. Consider any two vectors $\theta,\theta^{\prime}\in\mathbb{R}^{n}$ and the corresponding NEs $x(\theta),x(\theta^{\prime})\in X$, it follows from (11) that
$-G(x(\theta),\theta)\in\mathrm{N}_{X}(x(\theta)),\ -G(x(\theta^{\prime}),\theta^{\prime})\in\mathrm{N}_{X}(x(\theta^{\prime})),$
which immediately leads to
$\langle-G(x(\theta),\theta),x(\theta^{\prime})-x(\theta)\rangle\leq 0,\ \langle-G(x(\theta^{\prime}),\theta^{\prime}),x(\theta)-x(\theta^{\prime})\rangle\leq 0.$
From the above two inequalities, one has
$\langle-G(x(\theta),\theta),x(\theta^{\prime})-x(\theta)\rangle\leq\langle-G(x(\theta^{\prime}),\theta^{\prime}),x(\theta^{\prime})-x(\theta)\rangle,$
and hence, by Assumption 2-i) and 2-ii), it can be derived that
$\mu\|x(\theta^{\prime})-x(\theta)\|^{2}$ $\leq\langle G(x(\theta^{\prime}),\theta)-G(x(\theta),\theta),x(\theta^{\prime})-x(\theta)\rangle$
$\leq l_{\theta}\|\theta-\theta^{\prime}\|\|x(\theta^{\prime})-x(\theta)\|,$
Therefore, it has
$\|x(\theta^{\prime})-x(\theta)\|\leq\frac{l_{\theta}}{\mu}\|\theta-\theta^{\prime}\|.$ (34)
Subsequently, for any $\theta,\theta^{\prime}\in\mathbb{R}^{n}$, by applying (6), Assumption 1-iv), and (34), one can obtain
$|f_{i}(x(\theta),\theta)-f_{i}(x(\theta^{\prime}),\theta^{\prime})|$
$\leq|f_{i}(x(\theta),\theta)-f_{i}(x(\theta^{\prime}),\theta)|+|f_{i}(x(\theta^{\prime}),\theta)-f_{i}(x(\theta^{\prime}),\theta^{\prime})|$
$\leq(L_{x}\frac{l_{\theta}}{\mu}+L_{\theta})\|\theta-\theta^{\prime}\|.$
Consequently, the assertion (12) holds for $L_{F}=L_{x}\frac{l_{\theta}}{\mu}+L_{\theta}$. ∎
Appendix B Proof of Lemma 6
Proof.
Consider that $\theta_{k}$ is adapted to $\mathscr{F}_{k}$ and $u_{k}$ is independent of $\mathscr{F}_{k}$, in light of (15) in Lemma 2, one has
$\mathrm{E}[\hat{\nabla}F(\theta_{k})+\nabla\hat{\chi}_{\Theta}(\theta_{k})|\mathscr{F}_{k}]$
$=\mathrm{E}[\frac{n}{\xi}F(x(\theta_{k}+\xi u_{k}),\theta_{k}+\xi u_{k})u_{k}]+\nabla\hat{\chi}_{\Theta}(\theta_{k})$
$=\nabla\hat{F}(\theta_{k})+\nabla\hat{\chi}_{\Theta}(\theta_{k})$
$=\nabla\hat{\mathbf{F}}(\theta_{k}).$
It follows from the definitions of ^$\hat{\nabla}_{\varepsilon_{k}}F_{i}(\theta_{k})$ and ^$\hat{\nabla}F_{i}(\theta_{k})$, together with the triangle inequality and (6) that
^$\|\hat{\nabla}_{\varepsilon_{k}}F_{i}(\theta_{k})-\hat{\nabla}F_{i}(\theta_{k})\|$ $\leq\frac{n}{\xi}L_{x}\|x_{\varepsilon_{k}}(\theta_{k}+\xi u_{k})-x(\theta_{k}+\xi u_{k})\|$
$\quad+\frac{n}{\xi}L_{x}\|x_{\varepsilon_{k}}(\theta_{k})-x(\theta_{k})\|.$
Hence, it obtains that
^$\|\hat{\nabla}_{\varepsilon_{k}}F_{i}(\theta_{k})-\hat{\nabla}F_{i}(\theta_{k})\|^{2}\leq 4\frac{n^{2}}{\xi^{2}}L_{x}^{2}\varepsilon_{k}.$ (35)
Utilizing the $c_{r}$-inequality $\|\sum_{i=1}^{s}a_{i}\|^{2}\leq s\sum_{i=1}^{s}\|a_{i}\|^{2}$, it has
$\mathrm{E}[\|\hat{\nabla}_{\varepsilon_{k}}F(\theta_{k})-\hat{\nabla}F(\theta_{k})\|^{2}|\mathscr{F}_{k}]$
$\leq N\sum_{i=1}^{N}\mathrm{E}[\|\hat{\nabla}_{\varepsilon_{k}}F_{i}(\theta_{k})-\hat{\nabla}F_{i}(\theta_{k})\|^{2}|\mathscr{F}_{k}]$
which yields (30) by combining with (35).
As for (31), note that
$\mathrm{E}[\|\hat{\nabla}F(\theta_{k})+\nabla\hat{\chi}_{\Theta}(\theta_{k})-\nabla\hat{\mathbf{F}}(\theta_{k})\|^{2}|\mathscr{F}_{k}]$
$=\mathrm{E}[\|\hat{\nabla}F(\theta_{k})\|^{2}|\mathscr{F}_{k}]+\|\nabla\hat{F}(\theta_{k})\|^{2}-2\mathrm{E}[\hat{\nabla}F(\theta_{k})^{\textit{T}}\nabla\hat{F}(\theta_{k})|\mathscr{F}_{k}]$
$=\mathrm{E}[\|\hat{\nabla}F(\theta_{k})\|^{2}|\mathscr{F}_{k}]-\|\nabla\hat{F}(\theta_{k})\|^{2},$
where the last equality holds due to
$\mathrm{E}[\hat{\nabla}F(\theta_{k})^{\textit{T}}\nabla\hat{F}(\theta_{k})|\mathscr{F}_{k}]$ $=\mathrm{E}[(\hat{\nabla}F(\theta_{k}))^{\textit{T}}|\mathscr{F}_{k}]\nabla\hat{F}(\theta_{k})$
$=\|\nabla\hat{F}(\theta_{k})\|^{2}.$
Recall from Lemma 1 that $F(\theta)$ is $NL_{F}$-Lipschitz continuous. Then (31) holds by (16) in Lemma 2. Therefore, the proof is completed. ∎
Appendix C Proof of Corollary 2
Proof.
Applying the $\frac{nNL_{F}+1}{\xi}$-Lipschitz continuity of $\nabla\hat{\mathbf{F}}$ and substituting ^$\hat{\nabla}_{\varepsilon_{k}}F(\theta_{k})$ in (24) by ^$\hat{\nabla}F(\theta_{k})$, one can obtain that
$\hat{\mathbf{F}}(\theta_{k+1})$ $\leq\hat{\mathbf{F}}(\theta_{k})-\alpha\nabla\hat{\mathbf{F}}(\theta_{k})^{\textit{T}}(\hat{\nabla}F(\theta_{k})+\nabla\hat{\chi}_{\Theta}(\theta_{k}))$
$\quad+\frac{nNL_{F}+1}{2\xi}\alpha^{2}\|\hat{\nabla}F(\theta_{k})+\nabla\hat{\chi}_{\Theta}(\theta_{k})\|^{2}.$ (36)
Taking conditional expectation with respect to $\mathscr{F}_{k}$ on both sides of (C) and relying on (29) in Lemma 6, one has
$\mathrm{E}[\hat{\mathbf{F}}(\theta_{k+1})|\mathscr{F}_{k}]$
$\leq\hat{\mathbf{F}}(\theta_{k})-\alpha\|\nabla\hat{\mathbf{F}}(\theta_{k})\|^{2}+\frac{nNL_{F}+1}{\xi}\alpha^{2}\|\nabla\hat{\mathbf{F}}(\theta_{k})\|^{2}$
$\quad+\frac{nNL_{F}+1}{\xi}\alpha^{2}\mathrm{E}[\|\hat{\nabla}F(\theta_{k})+\nabla\hat{\chi}_{\Theta}(\theta_{k})-\nabla\hat{\mathbf{F}}(\theta_{k})\|^{2}|\mathscr{F}_{k}].$
The rest of the proof follows the same analysis as Theorem 1. ∎